# Kinematics of the Sagittarius Dwarf Spheroidal core: A 5D Analysis for a 6D Methodology with Gaia DR3

### v1

@@ 5 Conclusion
We anticipate that analyses along the lines that we consider, which include real distance measurements for stars, will reduce the uncertainties on measurements of the log-slope of the density profile and the velocity anisotropy in a similar fashion to our mock distances case. We predict that uncertainties on velocity anisotropy and log-slope can be reduced by an order of magnitude and a half order of magnitude respectively. These measured quantities can then be compared to analogues of Sgr in simulations to test for consistency of the $\Lambda$CDM model.
@@ 6 Acknowledegments
LS acknowledges support from DOE Grant desc0010813. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.

### v2

@@ 5 Conclusion
We anticipate that analyses along the lines that we consider, which include real distance measurements for stars, will reduce the uncertainties on measurements of the log-slope of the density profile and the velocity anisotropy in a similar fashion to our mock distances case. We predict that uncertainties on velocity anisotropy and log-slope can be reduced by an order of magnitude and a half order of magnitude respectively. These measured quantities can then be compared to analogues of Sgr in simulations to test for consistency of the $\Lambda$CDM model. Improved uncertainties can also be used to better constrain a dark matter annihilation or decay signal Evans et al. (2023).
We reiterate that in our analysis, we assume that the core of Sgr is in dynamical equilibrium, and that the presence of tidal disruption in the region of our studied sample may be neglected. Of course, we do not have definitive proof that this is the case. However, there is good motivation that our assumptions are at least a good approximation, given the following observations/theoretical results. First, in the sample that we identify, we do not detect statistically significant presence of streaming motion in the velocities in any of the coordinate directions we consider. Second, the observed velocity distribution is nearly gaussian in each of the coordinate directions, indicative of a system in near equilibrium. Third, from a theoretical perspective, the true dark matter mass for Sgr analogues in simulations has been shown to be well-determined assuming that the core of the system is in equilibrium and may be modeled using the jeans equations (Wang et al., 2022b; Tchiorniy & Genina, 2025). For these reasons, we believe that equilibrium is a good approximation to the system for the sample that we consider. We also make the assumption of sphericity, which is known to not be the case for Sagittarius. Previous studies have found $\sim 30$% differences between spherical and non-spherical jeans models (Bonnivard et al., 2015; Genina et al., 2018; Vitral et al., 2024).
@@ 6 Acknowledegments
The authors would like to thank Matthew Walker for his coding contributions, as well as Oz and Mac for their support. LS acknowledges support from DOE Grant desc0010813. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.
7 Data Availability
The data used in this article are publicly available from the European Space Agency (ESA) mission Gaia and can be accessed through the Gaia Archive. The resulting table of Sagittarius candidates used in this analysis is included as additional material in the file Sgr_candidates.txt.
@@ Appendix A Gaia Query
Appendix B Additional Analyses
Gaia proper motion errors are known to be underestimated, since they do not include a component for systematics (Lindegren et al., 2021). Vasiliev & Baumgardt (2021) quantify this effect with a correction factor added in quadrature. Taking a systematics correction for proper motion as suggested in Vasiliev & Baumgardt (2021), we perform a RAD+TAN Jeans analysis with both $\Delta_{D}=1$ kpc, $\Delta_{D}=0.1$kpc. Figure 10 compares the resulting posteriors to their counterpart without a systematics correction. No significant change is observed when the systematic error correction is taken into account.
Figure 10: Posteriors for a combined radial and tangential Jeans analysis. Brown and dark purple contours correspond to an analysis with no systematic error corrections and $\Delta_{D}=0.1$kpc, $\Delta_{D}=1$kpc respectively. Orange and pink contours correspond to an analysis with the systematic error correction for $\Delta_{D}=0.1$kpc, $\Delta_{D}=1$kpc respectively. The parameters are velocity anisotropy in the form $\log_{10}(1-\beta_{a})$, mass contained within the half light radius $\log_{10}(M_{HL}/M_{\odot})$, concentration $\log_{10}(c)$, and log slope at the half light radius LogSlope($r_{HL}$).